2. Consumer Price Index and Real Price Index: Over a four-year period, the prices of mangoes PM (in dollars) and the Consumer Price Index (CPI) are given below: Year CPI Price of Mangoes (PM) 1991 : cpi=94 , p=14 ; 1992: cpi=100 , p=15 ; 1993: cpi= 105 , p=16 ; 1994: cpi= 115 , p=17

2.1.Which year is the base year for CPI? Use this same base year for mangoes. 2.2.Calculate the nominal price index for mangoes. 2.3.Calculate the real price index for mangoes. 2.4.Have mangoes been more or less expensive compared to the typical goods consumers buy as measured by the CPI? Explain
Answer #1

1) 1992 is the the base year for CPI because CPI equals 100

2)

Year

CPI

Price

Nominal Price

Nominal PI - Mango

1991

94

14

13.16

93.33

1992

100

15

15.00

100.00

1993

105

16

16.80

106.67

1994

115

17

19.55

113.33

Working:

Year

CPI

Price

CPI current year/ CPI base year *100

Price/ Price base year

1991

94

14

14*94/100 = 13.16

14/15*100 = 93.33

1992

100

15

15*100/100 = 16

15/15*100 = 100

1993

105

16

16*105/100 = 16.8

16/15*100 = 106.67

1994

115

17

17*115/100 = 19.55

17/15*100 = 113.33

 

3)

Year

CPI

Price

Real PI

1991

94

14

99.28

1992

100

15

100.00

1993

105

16

101.59

1994

115

17

98.55

Working:

Year

CPI

Price

Real PI

1991

94

14

93.33*100/94 = 99.28

1992

100

15

100/100*100 = 100

1993

105

16

106.67 * 100/105 = 101.59

1994

115

17

113.33 * 100 /115 = 98.55

 

4) The abobe figure reglects that the average CPI exceeds the average real price index. We make a comparison of the the cost of living with the cost of mangos